OpenStudy (reeh2):

what do you need to do first to solve the square root of x plus the square root of x - 5 = 5

OpenStudy (anonymous):

So your equation is: \[\sqrt{x} + \sqrt{x - 5} = 5\]

OpenStudy (imstuck):

You need to square both sides. The first operation is quite tedious and kind of difficult, but it will eliminate the radicals.

OpenStudy (anonymous):

Or, is it: \[\sqrt{x} + \sqrt{x} - 5 = 5\]

OpenStudy (anonymous):

Wait a mo, it might not be the first equation. It depends on which one it is.

OpenStudy (imstuck):

You are correct; I just went with what I thought it was in my own mind! So...which one is it?!

OpenStudy (anonymous):

If it's: \[\sqrt{x} + \sqrt{x - 5} = 5\] Then square both sides like so: \[x + x - 5 = 25\] Then you can solve for x. If it's: \[\sqrt{x} + \sqrt{x} - 5 = 5\] Then add 5 to both sides to get x by itself. \[\sqrt{x} + \sqrt{x} = 10\] Then you can square both sides and solve for x.

OpenStudy (agent0smith):

@jcpd910 your first step doesn't work that way. \[\large (a+b)^2 \ne a^2 + b^2\] You'll probably have to square both sides twice.

OpenStudy (imstuck):

Your first step up above is incorrect, jcpd910. You do not end up with that at all.

OpenStudy (anonymous):

Woops. o_o